# Dynamical thermalization, Rayleigh-Jeans condensate, vortexes and wave collapse in quantum chaos fibers and fluid of light

### v2

@@ XI Discussion and conclusion
Appendix
@@
Appendix Figure A1: Color map of condensate probability $\rho_{1}$ at corresponding system temperature $T$ and number of system eigenmodes $N$ (left panel); right panel shows the same $\rho_{1}$ as in left panel but in the plane $(N,\epsilon_{0})$. Results are obtained in the frame of BE thermal distribution (1), here $\Delta=1$.
Appendix Figure A2: Dependence $S(t)$ from Figure 14 is shown here for an initial short time interval $0\leq t\leq 15$; averaging is done on a small time window $\delta t=0.1$ depicting initial oscillations between states $m=1$ and $m=2$ for initial state $\psi(t=0)=(\psi_{1}+\psi_{2})/\sqrt{2}$ at total energy $E=16.2$.
@@
Appendix Figure A4: Dependence of ground state probability $\rho_{g}(t)$ on time; same parameters as Figure 21 at $\beta=-10$ but the initial state is $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=5$ (left panel) and $m=7$ (right panel); blue/green curves are the probability in the linear/instanteneous ground state as in Figure 21 with the same dashed lines.
Appendix Figure A5: Same as Figure 22 at $\beta=-10$ and initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=5$ (top left panel, blue color) and $m=7$ (top right panel, green color); data are averaed over the range $200\leq t\leq 400$; dashed curves are as in Figure 21. Top panels show averaged probabilities $<\rho_{m}>$ vs. energies $E_{m}$ and corresponding bottom panels show entropy dependence on long/short times in right/left panels.

### v3

@@ XI Discussion
Acknowledgments: We thank S.Babin, D.Kharenko and E.Podivilov (IAE RAS Novosibirsk), K.M.Frahm (LPT Toulouse), and N.Pavloff (LPTMS Orsay) for useful discussions.
The authors acknowledge support from the grants ANR France project OCTAVES (ANR-21-CE47-0007), NANOX $N^{\circ}$ ANR-17-EURE-0009 in the framework of the Programme Investissements d’Avenir (project MTDINA), MARS (ANR-20-CE92-0041).
Appendix A
@@
Appendix Figure A1: Color map of condensate probability $\rho_{1}$ at corresponding system temperature $T$ and number of system eigenmodes NN (left panel); right panel shows the same $\rho_{1}$ as in left panel but in the plane $(N,\epsilon_{0})$. Results are obtained in the frame of BE thermal distribution (1), here $\Delta=1$.
Appendix Figure A2: Dependence of quantum von Neumann entropy $S_{q}(t)$ from Figure 14 is shown here for an initial short time interval $0\leq t\leq 15$; averaging is done on a small time window $\delta t=0.1$ depicting initial oscillations between states $m=1$ and $m=2$ for initial state $\psi(t=0)=(\psi_{1}+\psi_{2})/\sqrt{2}$ at total energy $E=16.2$.
Appendix Figure A3: Time dependence of the classical Boltzmann entropy $S_{B}(t)=\sum_{m}\ln\rho_{m}(t)$. The entropy is defined as $S_{B}=\sum_{m=1}^{m_{max}}\ln(\rho_{m})$, time average is done over a window of $\delta t=5$. The curves correspond to initial states $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ with $m$ values given in the panels; system parameters are the same as in Figure 14. The blue, green, and red curves show the cases with $\beta=10$ for $m=26,16,11$, respectively. The violet curve represents the case with $m=11$ and $\beta=-10$. The left panel uses a cutoff of $m_{max}=75$, while the right panel uses $m_{max}=150$. For each color, the semi-transparent curve shows the evolution of the entropy, while the solid, opaque curve represents its mean value, smoothed over a time window of $\delta t=5$.
@@
Appendix Figure A5: Dependence of ground state probability $\rho_{g}(t)$ on time; same parameters as Figure 21 at $\beta=-10$ but the initial state is $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=5$ (left panel) and $m=7$ (right panel); blue/green curves are the probability in the linear/instantaneous ground state as in Figure 21 with the same dashed lines.
Appendix Figure A6: Same as Figure 22 at $\beta=-10$ and initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=5$ (top left panel, blue color) and $m=7$ (top right panel, green color); data are averaged over the range $200\leq t\leq 400$; dashed curves are as in Figure 21. Top panels show averaged probabilities $\rho_{m}>$ vs. energies $E_{m}$ and corresponding bottom panels show the quantum von Neumann entropy dependence $S_{q}(t)$ on long/short times in right/left panels.